OpenStudy (anonymous):

What is the following number in decimal form? 1.75 × 10^−9 =

OpenStudy (anonymous):

@igreen ITS REVERSED!

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

I dont know what to do if its a negative

OpenStudy (igreen):

Yeah, you move the decimal place to the left instead of the right.

OpenStudy (anonymous):

10^-9 :/

OpenStudy (igreen):

\(2.25 \times 10^{-4} \rightarrow 0.000225\) Move the decimal place 4 times to the right.

OpenStudy (anonymous):

1.75 x 10^-9

OpenStudy (igreen):

*left

OpenStudy (igreen):

So for here we move it 9 places to the left.

OpenStudy (paxpolaris):

\[\large 1.75 \times 10^{−1} = 0.175\]\[\large 1.75 \times 10^{−2} = 0.0175\]\[\large 1.75 \times 10^{−3} = 0.00175\]\[\large 1.75 \times 10^{−4} = 0.000175\]

OpenStudy (anonymous):

i didnt know that!

OpenStudy (anonymous):

so it would be 1.00000000075

OpenStudy (anonymous):

\[1.75*10^{-9}=\frac{ 1.75 }{ 10^9 }=\frac{ 1.75 }{ 1,000,000,000 }=?\]

OpenStudy (paxpolaris):

decimal point .... followed by 8 zeroes ... followed by 175

OpenStudy (igreen):

Um, no, you just added more zeroes..you move the DECIMAL place over and add 0 before the 1.

OpenStudy (anonymous):

0.000000000175

OpenStudy (paxpolaris):

0.000 000 00 175

OpenStudy (paxpolaris):

for 10^-9 you will have 8 zeroes after the decimal

OpenStudy (anonymous):

\[\frac{ 1.75 }{ 1,000,000,000 }=\frac{ 175 }{ 100,000,000,000 }=?\]

OpenStudy (igreen):

Yeah, it's not 9 zeroes, because moving it one over gives you 0.175, not 0.0175